OpenStudy (anonymous):

1 cubic feet (1ft^3) of air is present in a piston cylinder arrangement. the axis of the cylinder is oriented in the same direction as gravity, with a 100 lbm, 1ft-dia piston on top of the air. The air surrounding the piston is at 14.7 psi. what is the absolute and gauge (gage) pressure of the air in the cylinder

OpenStudy (anonymous):

hello can any body help me

OpenStudy (ghazi):

do you know formula of gauge pressure

OpenStudy (anonymous):

P gauege = Pabs-Patm

OpenStudy (ghazi):

pressure = force per unit area....you have the diameter you can find pressure....to find force use F= mass * gravity

OpenStudy (anonymous):

but how can we use the mass of cylinder to find the pressure of the gas

OpenStudy (mathmate):

The piston diameter is 1 ft. The mass is 1 lb-mass. So the additional pressure (above atm) is \[ gauge\ pressure \ = \ mg/A\ = \ \frac{100\ lb.mass * 32.2 \frac{ft}{s^2} }{\frac{1}{4}\pi 1^2} \]

OpenStudy (mathmate):